Pregnancy Length in a Population of New Mothers 282 <x< 293 = 269 = 11 х 239 282 293 Pregnancy length (in days)
find the probability that the member selected at random is from the shaded area of the graph (round to four decimal places as needed)

please solve
Answer #1

Step 1: Sketch the curve. The probability that 282 < X < 293 is equal to the blue area under the curve. 0.045 0.040 0.035 0.0

Step 2: Since u = 269 and o=11 we have: 282 – 269 X-u P ( 282 < X < 293 ) = P ( 282 – 269 < X – u < 293 – 269 ) =P 293 – 269